OpenStudy (anonymous):

If a 7% saline solution and a 4% saline solution are mixed to make 500 milliliters of a 5% saline solution, how much of each solution, to the nearest milliliter, is needed? A. 167 milliliters of 7% solution and 333 milliliters of 4% solution B. 110 milliliters of the 10% solution and 210 milliliters of 4% solution C. 155 milliliters of the 10% solution and 300 milliliters of 4% solution D. 170 milliliters of the 10% solution and 340 milliliters of 4% solution

OpenStudy (anonymous):

Alg. 2?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

lol, same.

OpenStudy (anonymous):

FLVS?

OpenStudy (anonymous):

Or what?

OpenStudy (anonymous):

what

OpenStudy (anonymous):

NVM

OpenStudy (anonymous):

I'm not sure.

OpenStudy (anonymous):

I haven't gotten up to this.

OpenStudy (anonymous):

Either that or I forgot it.

OpenStudy (anonymous):

Let x be the number of milliliters of the 7% saline solution and y be the number of millilitres of the 4% saline solution. They need to add up to 500 milliliters total, so x + y = 500. The amount of salt in the x milliliter 7% solution is 0.07x milliliters. The amount of salt in the y millilitre 4% solution is 0.04y milliliters. Together, they need to add up to 5% of the total 500 milliliters, which is 0.05(500) = 25 milliliters, so 0.07x + 0.04y = 25. Solve by substitution: x + y = 500 x = 500 - y 0.07x + 0.04y = 25 (substitute 500 - y for x) 0.07(500 - y) + 0.04y = 25 35 - 0.07y + 0.04y = 25 -0.03y + 35 = 25 -0.03y = -10 y = 333.333... y = about 333 x = 500 - y = 500 - 333 = 167 get it?

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Okay them

OpenStudy (anonymous):

Then*